### v1

# Towards Fixing Panic Bugs for Real-world Rust Programs

Abstract.
The Rust programming language has garnered significant attention due to its robust safety features and memory management capabilities. Despite its guaranteed memory safety, Rust programs still suffer from runtime errors that are unmanageable, i.e., panic errors. Notably, over half of the bugs in rustc, Rust’s own compiler, are attributable to crash stemming from panic errors. However, understanding root causes and resolving these panics often requires substantial effort due to the limited information provided, and the stack backtrace could be intricate, often omitting the actual fault locations. Although numerous automated program repair techniques exist, we observe that the prevailing fix patterns do not readily apply to Rust programs due to natural differences in language mechanisms.
To tackle the above challenges, this paper introduces a systematic study aimed at fixing Rust panic bugs. We commence by assembling a dataset, namely Panic4R, which includes 102 real panic bugs and their fixes from the top 500 most downloaded open-source crates. By analyzing Rust’s implementation, we identify Rust-specific patterns for fixing panic bugs, which can aid in understanding and providing guidance for generating patches. Finally, we design and implement the first automated fixing tool, PanicKiller, for Rust panic bugs, which effectively generates correct patches on the real-world large-scale dataset, and has already assisted in the resolution of 28 panic bugs in open-source projects. Each resolved issue has been validated by the developers and merged into the respective codebases.
Rust, automated program repair, fault localization
1. Introduction
As a statically typed programming language, Rust has gained popularity for its well-known memory safety guarantees and high performance. Recently, the White House Office of the National Cyber Director also emphasized the necessity of using programming languages that have fewer memory safety vulnerabilities (WhiteHou52:online, ), and nominating Rust as an example of a memory-safe programming language. The foundational principles of Rust, including ownership, borrowing, and lifetimes, enable developers to implement secure and efficient programs. Rust’s emphasis on zero-cost abstractions and fearless concurrency has significantly contributed to its popularity in systems programming (bugden2022rust, ; jung2021safe, ; klabnik2023rust, ; jung2020understanding, ). This focus has led to an increase in the development of widely recognized software projects written in Rust (Servothe7:online, ; RedoxYou24:online, ; TiKVTiKV64:online, ; StratisS70:online, ; Cloudfla15:online, ).
Although Rust boasts security features and avoids common bugs such as null pointer dereference, uninitialized variables, and data races, it is still prone to panic errors, a Rust-specific crash type. The severity of the panic is evident from its typical consequences—program crashes or terminations that may lead to the improper handling of resources, such as unclosed file descriptors or network connections (Zheng2023A, ). Unlike Java’s structured exception handling framework that manages routine errors, Rust’s panic mechanism is designed for unrecoverable situations, significantly impacting program stability. Moreover, the Rust compiler, rustc, written in Rust, also exhibits vulnerabilities to panic bugs. Over half of the issues in Rust language’s official GitHub repository are categorized as Internal Compiler Errors (ICE) (Issues·r78:online, ), primarily caused by panic bugs in rustc. Therefore, understanding and resolving panic bugs in Rust is crucial for ensuring the reliability and stability of Rust programs.
However, fixing panic bugs can be a tedious and challenging task for Rust developers. On the one hand, the stack expansion information output by the compiler after a panic error is triggered could be complex and hard to understand. Regarding dozens of lines of stack backtrace, it is challenging to understand the root cause, localize the bug, and fix the program. Due to the unique memory management model and lifetime rules of Rust, the learning curve becomes particularly steep (Crichton2021The, ; crichton2024profiling, ; Qin2020Understanding, ). On the other hand, the infrastructure and toolchain supporting the Rust language are not as mature as other languages. Most of the existing code fix datasets or automated program repair (APR) tools are designed for other languages, such as Java and C/C++, but they could be difficult to migrate to Rust programs because of the differences in language mechanisms, which may result in the repaired program violating ownership rules or even failing to compile.
Recently, a few program repairing tools have been proposed for Rust, yet they are insufficient to address the most severe panic-related bugs. Rust-lancet (yang2024lancet, ) was developed to tackle bugs related to violations of ownership rules through three specific strategies. However, these strategies are tailored exclusively to Rust’s ownership rules, which are verified prior to runtime, making them unsuitable for addressing panic bugs. Similarly, Ruxanne (robati2024patterns, ) and RustAssistant (deligiannis2023rustassistant, ) focus on common compilation issues, such as incorrect data types, but these patterns do not effectively mitigate panic bugs.
To overcome the aforementioned challenges and fill the program understanding gap, in this paper, we design and implement the first infrastructure aiming at automatically fixing panic bugs for real-world Rust programs. We have constructed a dataset Panic4R, containing more than $100$ panic bug instances and corresponding fix patches, derived from open source projects in the ecosystem. Referring to the Rust implementation code, we further perform fix pattern mining with Rust-specific syntactic features to provide a reference for the community to better understand panic bugs and fixing strategies. Further, we implement an automated fixing tool for panic bugs, namely PanicKiller, which first applies dependency analysis for cross-file level error localization, then combines it with semantic information of the reported errors for fix pattern matching, and finally outputs sorted patches with scores and descriptions.
To evaluate the effectiveness of PanicKiller, we conducted extensive experiments on Panic4R. For fault localization, PanicKiller has achieved high accuracy at different granularities. We also compared PanicKiller to the LLM-based ChatGPT-4.0, employing both single and multi-round conversations as baselines. Results demonstrate that PanicKiller outperforms ChatGPT-4.0 in fault localization and patch generation, highlighting its practicality and reliability. Moreover, PanicKiller has effectively resolved issues in open-source Rust projects, with 28 panic bug fixes validated and merged by developers. In summary, the contributions of this work are as follows:
  • Dataset. We have constructed the first public dataset for Rust panic bugs, named Panic4R, which comprises 102 real bugs and their corresponding fixes from PR records of the top 500 most downloaded Rust crates. Each bug-fix pair is meticulously organized and has undergone thorough manual verification, facilitating future research.
  • Patterns. We mined a series of fix patterns for panic bugs compliant with Rust syntax. The potential application for these mined patterns could be developing automated repair tools, providing references for developers, serving as a dataset for fine-tuning LLMs, etc.
  • Tool. We introduced PanicKiller, an automated repair tool specifically for Rust panic bugs, designed to address issues in real-world and large-scale Rust programs. Our experiment results show that PanicKiller is more efficient than commercial LLM-based tools and has successfully resolved 28 open issues in Rust projects on GitHub.
The dataset, mined patterns, and the source code of PanicKiller can be found at https://sites.google.com/view/panickiller/home.
Figure 1. Overview of the proposed infrastructure.
2. Background and Motivation
Rust, a statically typed programming language, is gaining popularity for its well-known memory safety guarantees and high performance, making it notably distinct from C/C++ and JVM family of languages. The memory and thread safety is ensured through a specific mechanism known as the ownership system, which naturally prevents common errors like null pointer errors. Consider the code snippet presented in Listing 2, where line 2 declares a variable named name of type Option<String>. This type encapsulates a String object and employs Option, a Rust-specific data structure designed to handle potentially null values. It uses Some to indicate the presence of a specific value and None to represent the absence of any value, as demonstrated in line 3. In the subsequent match expression, the scenarios of Some and None for name are addressed separately. If name contains a value, this value is assigned to the variable n, which is then utilized to construct a new string. During this process, the ownership of n is transferred from name, rendering name subsequently null. As a result, in line 5, the ownership of name has been transferred to greeting via the match expression. Consequently, name is now None. Attempting to unwrap it in line 8 will trigger a compile-time error due to the ownership transfer. Also, as shown in line 9, calling unwrap() without checking the null value will result in a runtime error, i.e., panic error, and the program terminates because it cannot handle the exception.
{listing}
[h] {minted}[linenos=false, fontsize=, highlightcolor=gray!10, highlightlines=5,8,9]rust 1 fn main() 2 let name: Option¡String¿ = Some(String::from(”Alice”)); 3 let no_name: Option¡String¿ = None; 4 let greeting = match name 5 Some(n)=¿format!(”Hi, ”, n), //ownership transferred 6 None =¿ ”Hi, guest”.to_string(), 7 ; 8 // name.unwrap(); // compilation error 9 no_name.unwrap(); // PANIC 10 An example of Rust program.
Motivation. The distinctive memory checking mechanism and stringent ownership transfer rules in Rust present significant challenges for understanding Rust programs and debugging them. Besides, if a complex program encounters a panic bug at runtime, much effort is required to perform error localization and determine the root cause.
Although there have been many APR techniques practiced on Java/C/C++, due to the unique design mechanics of Rust, most of the existing repair patterns are not applicable to Rust programs. Specifically, Java uses null to represent a variable that points to no object in memory. While convenient for indicating uninitialized states, this design often leads to problems, with most of Java’s fix patterns addressing null pointer exceptions (liu2019tbar, ; koyuncu2019ifixr, ; kim2013par, ; liu2019avatar, ; durieux2017nullpointer, ). For example, the function in Listing 2 tries to return a null value while the expected return type is int. It would trigger a very common null pointer bug, and the existing APR tool generally adds an extra condition to check if the value is null, then return a random data that fits the type requirement.
{listing}
[h] {minted}[linenos=false, fontsize=, highlightcolor=gray!10, highlightlines=3,4,5]java public static int test() Integer value = null; // Not allowed in Rust + if (value == null) // Exisiting fix pattern + return 0; + return value; // NullPointerException A minimized Java program triggering null pointer exception and the corresponding fix patch.
However, conventional null pointer remediation strategies are largely ineffective in Rust due to the language’s rigorous handling of null values. As demonstrated in Listing 2, Rust does not employ a direct equivalent of null. Instead, it utilizes the Option<T> data type to explicitly manage scenarios where values might be absent. Data that could be null must be encapsulated within an Option<T>, effectively preventing null pointer errors inherent to Rust’s design. Consequently, traditional fix patterns that address null pointer issues do not apply to Rust programs. Besides, we notice that the C/C++’s bug fixing dataset, CVE-Fixes (fu2022vulrepair, ; chen2023vrrepair, ), covers a wide variety of memory bugs such as improper restriction of operations within the bounds of a memory buffer (CWE-119), out-of-bounds write (CWE-787), and null pointer dereference (CWE-476), etc. However, these specific bug types are inherently untriggerable in safe Rust programs due to the language’s stringent safety guarantees. Thus, the methods commonly applied in C/C++ contexts are not transferable to Rust programs.
3. Infrastructure
The workflow and architecture are illustrated in Figure 1. Initially, we gather real-world panic bugs and their corresponding patches from the ecosystem to construct a dataset, Panic4R, the specifics of which are detailed in Section 3.1. Subsequently, referring to the implementation of Rust’s compiler rustc, and its standard libraries, we identify panic-fix patterns, including abstract patches and natural language descriptions, as elaborated in Section 3.2. Finally, we develop an automated fixing tool, PanicKiller, utilizing the extracted patterns. PanicKiller conducts dependency analysis on source programs, organizes patch priorities, and generates hybrid fixing suggestions, further explained in Section 3.3.
3.1. Dataset Construction
To build a dataset containing real-world panic bugs and their patches, we follow the construction process of Defects4J (just2014defect4j, ), which is the most classic dataset in the APR task, and select the top $500$ downloaded crates from the public repository of Rust crates (cratesio6:online, ) for real bugs and fix patches extraction, so as to construct a real dataset. Specifically, we manually employ the following workflow to collect code and thus guarantee the quality and reliability of the dataset.
Table 1. Statistics of bugs and patches available in Panic4R.
Download
Ranking
# Bugs
LoC
Test LoC
Crates Involved
1-50 36 25,882 13.60 syn, rand, regex, aho-corasick, num-traits, clap, serde_json, strsim, time, idna, hashbrown, proc-macro2, smallvec
51-100 35 25,988 13.58 percent-encoding, chrono, uuid, textwrap, nom, tokio, hyper, futures, toml
101-150 6 26,516 23.83 httparse, object, rustc-demangle, rustls, form_urlencoded, gimli
151-200 3 19,173 16.67 reqwest, num-bignit, rayon
201-250 6 3,616 9.70 bumpalo, filetime, fixedbitset, phf, prost
251-300 3 25,572 10.33 pest, serde-yaml, libm
301-350 2 44,882 10.50 prettyplease, bytemuck
351-400 2 1,853 12.00 cargo_metadata, tinytemplate
401-450 5 5,444 25.00 tar, plotters, pretty_assertions, yansi
451-500 4 28,760 23.25 crossbeam, brotli-decompressor, indicatif, md5
Total 102 21,805 15.87 51
Table 2. Partial fix patterns derived from the Rust implementation code.
Fix Patterns Code Changes Interpretation Templates Example PRs
Insert Match Unwrapper
-   x = exp.unwrap();
+   x = match exp.unwrap() {
+     Some(_) => { exp.unwrap() }
+     _ => { return }
When unwrapping on [value], add match arms to [variable] after unwrapping to handle all possible circumstances to avoid panics caused by unwrapping on None/Invalid values.
serde_json (PR 757)
nom (PR 1032)
Reorder State Changer
//  polling operation
+  stmt1 //  advance statement
//  other statements
-  stmt1 //  state changer
Advance the statement [state changer] to avoid incorrect state resumption after asynchronous functions have completed.
futures (PR 2250)
opendal (PR 4013)
Delete Second Borrow
data.borrow() //  immutable borrow
-  data.borrow_mut()
Delete the second mutable borrow of [data] when there exists immutable borrow to avoid ownership violation panics.
StackOverflow (SO 1)
Mutate Error Handler
-  x.expect()    +  x.unwrap()
//  or unwrap_or_default/else
Replace the [original handler] with [new handler] to avoid panics caused by incorrect error handling like [original handler].
clap (PR 4480)
Mutate Binary Operator
-   a op b
1. Mutate to wrapping/saturating function:
+  a.wrapping/saturating_op(b)
2. Mutate to checked function:
+   a.checked_op(b).unwrap()
//  or unwrap_or_default/else
Replace basic arithmetic operations [operator] with safer operations [call name] to handle arithmetic [operator] overflow panics. Note that [explanation].
regex (PR 996)
chrono (PR 1294)
chrono (PR 1023)
chrono (PR 686)
Insert Range Checker
1. Check the index:
+   if index > arr.len() { return }
2. Check start/end of range if index is range:
+   if end > x.len() { return } or
+   if start >= x.len() { return }
3. Check whether start $>$ end if index is range:
+   if start > end { return }
Implement range checking for the [index] of indices [array name] to determine whether [condition], avoiding index out of bounds or exceed the boundary.
idna (PR 658)
idna (PR 655)
Mutate Index Expression
-  array[index1]    +  array[index2]
Mutate [index] in indices [array name], avoiding index out of bounds or exceeding the boundary.
textwrap (PR 391)
Mutate Condition
-  if cond1     +  if cond2 && cond1
//  add after if-let expression
Adjust conditions within if statements to check whether [condition].
idna (PR 865)
Insert Unsafe Block
//  necessary condition
-  exp1    +  unsafe { exp2 }
Insert an unsafe block when [precondition] is met to change the behaviour of [variable].
nom (PR 370)
Mutate Method Invocation
-  x.y([params])     +  x.z([params])
Replace the original call [call name] with another [new call name] with the same parameters.
hyper (PR 2410)
serde_json (PR 493)
Insert Call Invocation
-  x.y()    +  x.y().z()
Add new method call [call name] to [variable].
chrono (PR 1254)
nom (PR 1618)
(1) Identifying Real Fix Patches. For each target crate, we review the list of pull requests (PRs). If a closed PR’s title or description contains keywords like “fixing/repairing panic,” it is considered a potentially valid fix patch. We also assess PRs linked to issues with keywords such as “panic/thread panicked at…”. Then, we further examine the content of the code changes corresponding to PRs and their related PRs to ensure that they contain fixes for a panic bug. For example, PRs that merely add or remove test cases without modifying the crate’s source code are not considered relevant to our dataset. Similarly, changes to the API documentation for the crate are also outside the scope of our dataset construction. We only select closed PRs to ensure that they are either approved by developers or submitted by the developers themselves.
(2) Reproducing Real Bugs. To ensure the reproducibility of panic bugs in each crate, which requires specific versions, we download the source code for the corresponding commit IDs before and after the PR commit. The two versions of the target crate are regarded as $V_{bug}$ and $V_{fix}$, respectively. Then, we refer to the description in the PR or the code in the corresponding issue to minimize the test case $T_{test}$ that triggers the bug, i.e., calling the API of the target crate. We ensure that $T_{test}$ can trigger panic bugs and be compiled before and after switching versions, respectively.
(3) Isolating Real Bugs. In the target crate’s repository, each PR commit may contain more than one code change, such as adding other functional logic or modifying data structures for new features, but not as a fix patch for panic. To ensure the accuracy of the fix patches in the dataset, we manually verified code changes and kept only the patches used to fix the panic bug. If a PR contains multiple fixes for panic bugs, we also split it into multiple code-fix pairs. As a result, the patches for $V_{fix}$ versus $V_{bug}$ have no irrelevant content and are the precise changes necessary to fix the panic bug. To summarize, each of the real datasets we build contains two crate versions, $V_{bug}$ and $V_{fix}$, and a test case $T_{test}$ that triggers panic bugs on $V_{bug}$ and is compilable on $V_{fix}$. Table 1 shows the details of the real-world dataset Panic4R.
3.2. Fix Pattern Mining
To comprehensively identify the causes and fix patterns for panic bugs in Rust programs, we analyze and infer data from the Rust implementation code (rustlang24:online, ). In the Rust compiler source code, panic-related error messages are encapsulated within macros such as panic_const!(). These macros and their accompanying messages provide the diverse types of panics and their fundamental causes. We compile a summary of these categories and their descriptions in Table  3. Following this categorization, we examine specific scenarios within each category. Rust compiler developers mark each panic occurrence with the annotation # Panics, which details the causes and circumstances of the panic. These annotations are often paired with # Safety or # Examples, providing either fix strategies or bug-triggering examples. This annotated information allows us to delineate code-level fix patterns for each panic type.
For example, Listing 3.2 showcases a Rust source code snippet annotated to indicate a panic bug related to the ownership mechanism. The annotation includes an example where the panic is triggered by a double borrowing of a RefCell parameter, which is a Rust smart pointer facilitating internal mutability. To prevent such panics, the annotation suggests using the try_borrow() method as an alternative approach. By systematically reviewing all documented panic instances and their proposed resolutions in the Rust source code, we collect a comprehensive list of fix strategies.
{listing}
[h] {minted}[linenos=false, fontsize=]rust /// # Panics: Panics if the value is currently /// mutably borrowed. For a non-panicking variant, /// use [‘try_borrow‘] (#method.try_borrow).
/// # Examples: An example of panic: /// let c = RefCell::new(5); /// let m = c.borrow_mut(); /// let b = c.borrow(); // this causes a panic pub fn borrow(&self) -¿ Ref¡’_, T¿ … The implementation of the borrow() function in RefCell, including documentation about potential panics.
To further delineate fix patterns, we convert code modifications into Abstract Syntax Trees (ASTs). By omitting irrelevant tokens, such as variable names, we identify similar AST transformation structures that reveal potential fixes for each i dentified root cause. In total, we collected 19 fix patterns, of which a representative categorization and corresponding examples are shown in Table 2, and the complete list of patterns can be found on our website.
Table 3. Root causes derived from panic bugs in real world.
Root Causes Code Examples
Unwrap on None/Invalid value x.unwrap() // when x = None
Mixed borrowing
x.borrow() x.borrow_mut()
Async functions wrong resume
// wrong state change
Arithmetic overflow
i32::MAX + 1
Index out of bounds
array[len + 1..]
Invalid UTF-8 boundary
&s[1..] // when s = "™";
Division/Modulus by zero
a / b // when b = 0
Assertion failed
assert!(1 == 0)
Unreachable code
unreachable!()
Others
panic!("A panic!")
3.3. Automated Panic Bugs Fixing Tool
Based on the mined fix patterns, we design a pattern-based automated panic bugs fixing tool, namely PanicKiller. PanicKiller first locates suspicious expressions based on the backtrace information output by the compiler and the dependency flow analysis of the original program. Then, combined with the semantic information of the errors, PanicKiller generates a series of patches for each fault location. Finally, PanicKiller sorts all the patches with verification and matching scores calculation, and outputs the top-5 ranked patches along with the corresponding natural language interpretation.
3.3.1. Fault Localization
When a panic occurs, Rust initiates the unwinding process, meticulously tracing back up the stack to ensure the cleanup process. The error message is further supplemented with backtrace information, and based on it, PanicKiller initially extracts suspicious locations, including file paths, column and row numbers. Considering that the precise locations of bugs might not be explicitly disclosed in the backtrace (schroter2010stacktrace, ), prior research suggests that the actual buggy location is apt to exhibit structural resemblances to the expressions detailed in the backtrace. This indicates that the exact location could be involved in other expressions and is dependent on those specified within the backtrace.
To capture the dependency relationship, we construct a Code Element Dependency Graph (CEDG) by utilizing the High-level Intermediate Representation (HIR), which serves as one of the intermediary forms during Rust’s compilation process. For the assignment statements or expressions, the variables on the left-hand side values are dependent on the right-hand side values. When it comes to function invocations, there may exist dependencies in the parameters of functions, so we delve deeper into the function to uncover additional dependencies. Leveraging the transitivity of the dependency relationship, we iteratively construct the CEDG. Then, based on the constructed dependency relationship, we first define the confidence score of each localized element $e_{i}$ as follows:
(1) $\small Con_{i}=1-\frac{\min(Dist(\text{backtrace},e_{i}),\lambda)}{\lambda},\quad\lambda\geq 1\$
where $Dist(\text{backtrace},e_{i})$ denotes the shortest distance on the dependency graph between element $e_{i}$ and any of the elements in the backtrace. A constant $\lambda$ is used to normalize the confidence score thus ensuring it is scaled within a meaningful range. In our work, we set $\lambda=2$ according to an existing study (laura2014dep, ), which has proved that it has the optimal performance. It involves incorporating both the precise locations and the code elements dependent on those identified, ensuring a more comprehensive analysis.
In addition to the confidence score, the presence of suspicious files within the backtrace can also impact fault localization. On the one hand, files that frequently appear in the backtrace tend to be more critical to the execution path leading to the error, making them more likely to be associated with the fault. On the other hand, within an expanded backtrace, a shallower depth suggests a stronger connection to the error’s origin (wong2014stacktrace, ). Thus, PanicKiller computes the suspicion score $Loc_{i}$ for each location as follows:
(2) $\small Loc_{i}=N\times\left(1/D+Con_{i}\right)$
where $N$ denotes the number of times a particular file of $e_{i}$ occurs within the backtrace, and $D$ represents the depth of the location in the backtrace sequence. For each buggy source code and the corresponding backtrace, PanicKiller calculates the score for each element $e_{i}$ after constructing a dependency graph to determine the ranking of a suspect location.
3.3.2. Pattern-based Patch Generation
Based on the suspicious locations, PanicKiller iteratively attempts to match each location with the mined fix patterns, which are illustrated in Section 3.2. Considering that there may be multiple suspicious locations and matched patterns, resulting in multiple combinations of fixes, we sort them based on validation.
Patch validation. For the APR task, it is important to guarantee that the generated patches repair the bug as well as do not affect the original semantics. In our work, we incorporate cargo-test, a built-in tool in Rust’s package manager, to automate the execution of regression tests. After performing the validation, the patches would have several cases: (1) The panic is eliminated and all the test cases execute with the same result as before. Our goal is to generate these patches, which indicate the most likely correct fix. (2) The panic is eliminated but the execution results of some test cases are not consistent with the original program. This suggests that the patch may have introduced logical modifications that cause the semantic inconsistency. (3) The panic is not eliminated and we don’t evaluate the regression testing because it’s not a correct fix. By regression testing, we can effectively filter out patches that might introduce new errors or exhibit overfitting to test suites, thereby enhancing the reliability and robustness of fixing.
Patch prioritization. For each suspicious location, PanicKiller iteratively employs the mined fix patterns. If the corresponding AST has a structure that matches a specific pattern, it is applied to generate a patch; otherwise, it is discarded. Note that even after a pattern has been applied to an expression, additional patterns might be identified as one delves deeper into the AST. Consequently, a suspicious location may correspond to more than one resulting patch. For all the generated patches, PanicKiller combines the scores and validation results mentioned above to perform the patch ordering. Specifically, patch prioritization is guided by three factors: (1) the confidence score for fault localization $Con_{i}$, (2) the similarity score of the patch’s interpretation, the templates of which are shown in Table 2, to the error messages, and (3) the results from the regression test validation. PanicKiller ranks each patch based on the cumulative sum of the two scores. For patches that achieve identical scores, preference is given to those that successfully pass regression testing.
Finally, PanicKiller outputs a top-5 ranked list of mixed-form fixing recommendations. To address patch interpretability issues (huang2023survey, ), PanicKiller provides each code patch alongside its natural language explanation, as detailed in Table 2. This description ensures developers receive a clear understanding of the modifications and the patterns applied in each patch.
Figure 2. A panic bug-triggering example and its patches.
A running example. Figure 2 presents a code example that triggers a panic bug caused by an index out-of-bounds error. Assuming that PanicKiller has identified lines $1$ and $4$ as suspicious locations, with confidence scores of $0.7$ and $1$, respectively, as determined by Equation 2. Subsequently, PanicKiller iteratively applies each pattern to these locations. For line $1$, PanicKiller identifies a method call of unwrap(), thus the Insert Match Unwrapper pattern is applied, and patch 1 is depicted in Figure 2 (b). For line $4$, PanicKiller initially identifies a binary expression containing a subtraction operator, and the Mutate Binary Operator pattern is employed, as illustrated as patch 2 in Figure 2 (c). Further, PanicKiller analyzes the deeper structure and identifies an index expression, for which the Insert Range Checker pattern is applied to arr[num2]. Patch 3 is depicted in Figure 2 (d).
Table 4. Prioritization results for the three patches.
Patch Fix Pattern Score ($Con_{i}$+$Loc_{i}$) Validation Rank
Patch 1 Insert Match Unwrapper 1.1 (0.7 + 0.4) Test fail 3
Patch 2 Mutate Binary Operator 1.1 (1.0 + 0.1) Test pass 2
Patch 3 Insert Range Checker 1.7 (1.0 + 0.7) Test pass 1
The prioritization of patches is shown in Table 4. Patch 3 ranks first with the highest score of $1.7$, having passed the regression tests. Patch 2 is second because, although it scored the same as patch 1, it passed the regression tests while patch 1 failed. Consequently, PanicKiller outputs this ranked patch list with their interpretations, organized by scores and test results.
4. Experiment Design
In this section, we illustrate our research questions and the baseline approaches employed in our experiments. We have implemented PanicKiller with Rust, and captured Rust’s High-level Intermediate Representation (HIR) (hir:online, ), which is a special representation used in the Rust compiler, to construct the dependency graph of suspicious locations. In addition, we implement the visiting and modification of the AST nodes based on the AST manipulation library syn (syn:online, ), to generate the fix patches. All the experiments are constructed on a Ubuntu 20.04.1 LTS laptop with 11th Gen Intel(R) Core(TM) i5-11300H @3.10GHz CPU.
4.1. Research Questions
PanicKiller is designed to automatically fix Rust panic bugs. To this end, we empirically explore the following research questions (RQ):
RQ1 (Fault Localization): To what extent does PanicKiller provide reliable fault localization results for panic bugs fixing? In this RQ, we aim to evaluate the accuracy of our bug localization approach based on backtrace information and dependency analysis. We categorize the real data in Panic4R into small-scale datasets ($<20$K LoC) and large-scale datasets ($\geq 20$K LoC) based on the size of the source code. We present the accuracy of various localization approaches at top-1, top-3, and top-5 levels. The accuracy is calculated based on three granularity levels: files, statements, and expressions.
RQ2 (Fixing Effectiveness): How effective is PanicKiller for fixing panic bugs compared to state-of-the-art LLM-based techniques? To evaluate PanicKiller’s effectiveness, we first verified the accuracy of patches it generated on Panic4R. Additionally, we conducted comparative experiments to assess patch generation accuracy when correct error localization information was given to both PanicKiller and baseline methods. Referring to existing research on APR (qi2015plausible, ), we separately verify the proportion of fixes that are plausible, i.e., the panic bug is successfully fixed, but some of the regression test cases fail, and the proportion of correct fixes, i.e., the panic bug is successfully fixed, and the regression test cases all pass.
RQ3 (Efficiency): What is the fixing efficiency of PanicKiller compared with traditional manual repair methods? In this RQ, we measure the time required to apply PanicKiller for fixing each target, and compare it with the duration from when the issue was reported to the PR was merged. This comparison aims to assess the extent to which our proposed approach enhances the efficiency of bug repairing.
Table 5. The correctness of tools on real and synthetic datasets, respectively, with different granularities of error localization.
Tool Panic4R-Small (61) Panic4R-Large (41) Total (102)
File Stmt Expr File Stmt Expr File Stmt Expr
Top-1 panic 41 (67.2%) 30 (49.2%) 30 (49.2%) 22 (53.7%) 10 (24.4%) 10 (24.4%) 63 (61.8%) 40 (39.2%) 40 (39.2%)
random 33 (54.1%) 14 (23.0%) 13 (21.3%) 14 (34.1%) 7 (17.1%) 6 (14.6%) 47 (46.1%) 21 (20.6%) 19 (18.6%)
similarity 13 (21.3%) 0 (0%) 0 (0%) 7 (17.1%) 1 (2.4%) 0 (0%) 20 (19.6%) 1 (1.0%) 0 (0%)
GPT-4 42 (68.9%) 28 (45.9%) 19 (31.1%) 10 (24.4%) 4 (9.8%) 3 (7.3%) 52 (51.0%) 32 (31.4%) 22 (21.6%)
GPT4-multi 43 (70.5%) 28 (45.9%) 19 (31.1%) 10 (24.4%) 4 (9.8%) 3 (7.3%) 53 (52.0%) 32 (31.4%) 22 (21.6%)
PanicKiller 45 (73.8%) 33 (54.1%) 32 (52.5%) 22 (53.7%) 13 (31.7%) 13 (31.7%) 67 (65.7%) 46 (45.1%) 45 (44.1%)
Top-3 random 46 (75.4%) 29 (47.5%) 29 (47.5%) 17 (41.5%) 8 (19.5%) 8 (19.5%) 63 (61.8%) 37 (36.3%) 37 (36.3%)
similarity 21 (34.4%) 1 (1.6%) 1 (1.6%) 10 (24.4%) 1 (2.4%) 0 (0%) 31 (30.4%) 2 (2.0%) 1 (1.0%)
GPT-4 43 (70.5%) 29 (47.5%) 20 (32.8%) 12 (29.3%) 4 (9.8%) 3 (7.3%) 55 (53.9%) 33 (32.4%) 23 (22.5%)
GPT4-multi 44 (72.1%) 29 (47.5%) 20 (32.8%) 12 (29.3%) 4 (9.8%) 3 (7.3%) 56 (54.9%) 33 (32.4%) 23 (22.5%)
PanicKiller 49 (80.3%) 38 (62.3%) 37 (60.7%) 28 (68.3%) 19 (46.3%) 18 (43.9%) 77 (75.5%) 57 (55.9%) 55 (53.9%)
Top-5 random 48 (78.7%) 33 (54.1%) 33 (54.1%) 22 (53.7%) 13 (31.7%) 12 (29.3%) 70 (68.6%) 46 (45.1%) 45 (44.1%)
similarity 25 (41.0%) 1 (1.6%) 1 (1.6%) 13 (31.7%) 1 (2.4%) 0 (0%) 38 (37.3%) 2 (2.0%) 1 (1.0%)
GPT-4 46 (75.4%) 30 (49.2%) 21 (34.4%) 14 (34.1%) 4 (9.8%) 3 (7.3%) 60 (58.8%) 34 (33.3%) 24 (23.5%)
GPT4-multi 47 (77.0%) 30 (49.2%) 21 (34.4%) 14 (34.1%) 4 (9.8%) 3 (7.3%) 61 (59.8%) 34 (33.3%) 24 (23.5%)
PanicKiller 49 (80.3%) 38 (62.3%) 37 (60.7%) 29 (70.7%) 20 (48.8%) 19 (46.3%) 78 (76.5%) 58 (56.9%) 56 (54.9%)
RQ4 (Case Study): Can PanicKiller be applied to the real-world Rust projects and fix practical panic bugs? To evaluate the practical effectiveness, we apply PanicKiller to fix real-world issues in GitHub, which are not included in Panic4R. For open issues, we submit a PR with the fix patch generated by PanicKiller, and we track whether they are confirmed and merged by developers. For closed issues, we use TF-IDF (sparck1972statistical, ) to calculate the similarity between the generated patch with the official patch. If the score exceeds $0.75$, and the panic is fixed with the regression test passed, we consider they are similar. If the PR is accepted by project developers or if our generated patch is similar to the official patch, it could further demonstrate that PanicKiller can be employed to real-world application scenarios and provide fixing guidance for large-scale real-world projects.
4.2. Baseline Approaches
In bug localization, while existing techniques typically rely on additional natural language descriptions from bug reports or test suites (yoo2012regression, ), our tool, PanicKiller, solely utilizes the error message from the Rust compiler and the accompanying backtrace. The absence of extra data makes traditional techniques unsuitable as baselines for our approach. Furthermore, since these techniques are designed for languages like C, C++, Java, and Python, they are not tailored to handle panic errors in Rust, which have distinct error message characteristics. Consequently, we devised three baseline methods based on conventional localization approaches: (1) panic: Directly using the location indicated in the panic error message. (2) random: Randomly selecting a suspicious location from the backtrace. (3) similarity: Calculating textual similarity between the error message and each code location; the highest similarity score indicates the fault location.
To our knowledge, no existing tools specifically address bug localization or the automated repair of Rust panic bugs. We therefore benchmark against the state-of-the-art LLM-based technique, ChatGPT. Due to ChatGPT-3.5’s limitations in handling files, we utilize the advanced and commercial ChatGPT-4.0 as our primary baseline. For each target, we upload a zip archive containing the entire Rust project, the test case triggering the panic, and relevant panic information. We then use a consistent prompt template, as shown below:
I have uploaded a source code package of a Rust crate [crate-name], [crate-zip]. When using this crate with the following test case: [main.rs], it went panic, the panic information is as follows: [panic_info].
[Location Prompt] + Based on the locations, provide the top 5 fixing code to fix this panic. Don’t explain anything about code, just show me the suspicious locations and patches.
As for the [Location Prompt], because we verify the correctness of the generated patches with/without the correct error location, the template are specifically as follows:
  • Without Perfect Location: Please give me the top 5 suspicious fault locations in the crate, including their path, line number and column number.
  • With Perfect Location: The error location is: [perfect-location].
Because ChatGPT may fail to produce bug localization or repair patches as instructed, we employ a baseline involving multiple interactions with ChatGPT. Initially, we assess ChatGPT’s output for completeness and iterate the query process up to three times, correcting for any missing information. If, after three attempts, localization or repair details remain unattained, we classify the effort as a ChatGPT repair failure. The inquiry prompts for missing information are as follows:
  • Localization: Please provide me directly with the exact file path, code line and column numbers where the error was located.
  • Fixing: Based on the mislocalization information you provided, please provide me with the specific code to fix it.
5. Result Analysis
In this section, we present the experiment results on automatically fixing panic bugs and then analyze the effectiveness and efficiency of our proposed tool PanicKiller.
5.1. Answer to RQ1: Fault Localization
To comprehensively assess the impact of fault localization through various techniques, we employ three granularity levels for locating panics: file, statement, and expression levels. As illustrated in Table 5, the localization accuracy of PanicKiller at all three granularities is higher than the other baselines, which indicates the effectiveness of PanicKiller. In addition, all methods perform better on the small-scale dataset compared with the large-scale dataset. This result is as expected, because on large-scale projects, the stack unfolding information could be more complex and harder to analyze the dependency for fault localization.
From the results, the localization accuracy of ChatGPT4 is basically the same for single and multi-round conversations due to a limited understanding of an entire Rust project. Throughout our experiments, we have encountered ChatGPT failing to parse the file and only outputting some broad suggestions, which is not useful for practical fault localization. Interestingly, the effectiveness of the random method is relatively close to PanicKiller, especially on small-scale datasets and top-5 prediction results, which suggests that if the localization is in the stack-expanded list, random selection can be effective, but as soon as in-depth dependency analyses are required +on large-scale dataset?, the random method fails outright. The similarity-based selection techniques are the least effective in terms of fine-grained localization accuracy, implying that relying on natural language alone does not accurately capture the semantic information of complex practical code, resulting in low localization accuracy.
5.2. Answer to RQ2: Fixing Effectiveness
Because there is no existing APR tool for Rust panic bugs, we focus on comparison the fixing effectiveness with ChatGPT-4. The evaluation results are shown in Table 6. Overall, we can conclude that PanicKiller significantly surpasses the performance of automated repairs conducted by GPT-4. Notably, within the large-scale dataset, GPT-4 fails to generate any viable patches, the reason of which is closely tied to GPT-4’s challenges with fault localization, as elaborated in Section 5.1. Moreover, we find that LLM-based methods frequently overlooks the code context when generating patches, resulting in the generated program failing to compile instead. For example, it may attempt to resolve panics by altering a method’s return type directly or by incorporating new logic, which results in both compilation and semantic errors.
Table 6. The number of correct patches and correctness of different tools on Panic4R, respectively.
Tool Panic4R-Small (61) Panic4R-Large (41) Total (102)
Plausible Correct Plausible Correct Plausible Correct
Top-1 GPT-4 6 9.8% 6 9.8% 0 0% 0 0% 6 5.9% 6 5.9%
GPT4-multi 8 13.1% 8 13.1% 0 0% 0 0% 8 7.8% 8 7.8%
PanicKiller 29 47.5% 19 31.1% 10 24.4% 7 17.1% 39 38.2% 26 25.5%
Top-3 GPT-4 6 9.8% 6 9.8% 0 0% 0 0% 6 5.9% 6 5.9%
GPT4-multi 8 13.1% 8 13.1% 0 0% 0 0% 8 7.8% 8 7.8%
PanicKiller 32 52.5% 23 37.7% 13 31.7% 10 24.4% 45 41.1% 33 32.4%
Top-5 GPT-4 7 11.5% 7 11.5% 0 0% 0 0% 7 6.9% 7 6.9%
GPT4-multi 9 14.8% 9 14.8% 0 0% 0 0% 9 8.8% 9 8.8%
PanicKiller 35 57.4% 25 41.0% 14 34.1% 10 24.4% 49 48.0% 35 34.3%
As shown in Table 7, even with perfect location information, PanicKiller consistently outperforms GPT-4. This indicates that PanicKiller’s patch generation capability is superior to that of GPT-4, even after multiple iterations. Compared with the results in Table 6, PanicKiller only slightly improves its fixing effectiveness with perfect location information. While it does generate new plausible or correct patches, it sometimes fails to recreate some of its previous patches without perfect locations. This is because some patches were applied at related but different locations, such as within a method call hierarchy, and still effectively fixed the issue. However, the provided location may interfere with the selection of fix patterns. As for ChatGPT, when provided with exact fault locations, it can generate a few effective patches for large datasets, a stark improvement over its complete failure without precise locations. This further underscores the deficiency in GPT-4’s fault localization capabilities.
Table 7. The number of correct patches and correctness of different tools on Panic4R with perfect locations, respectively.
Tool Panic4R-Small (61) Panic4R-Large (41) Total (102)
Plausible Correct Plausible Correct Plausible Correct
Top-1 GPT-4 6 9.8% 4 6.6% 3 7.3% 3 7.3% 9 8.8% 7 6.9%
GPT4-multi 6 9.8% 4 6.6% 3 7.3% 3 7.3% 9 8.8% 7 6.9%
PanicKiller 33 54.1% 24 39.3% 10 24.4% 7 17.1% 43 42.2% 31 30.4%
Top-3 GPT-4 6 9.8% 5 8.2% 3 7.3% 3 7.3% 9 8.8% 8 7.8%
GPT4-multi 6 9.8% 5 8.2% 3 7.3% 3 7.3% 9 8.8% 8 7.8%
PanicKiller 36 59.0% 27 44.3% 13 31.7% 11 26.8% 49 48.0% 38 37.3%
Top-5 GPT-4 6 9.8% 5 8.2% 4 9.8% 4 9.8% 10 9.8% 9 8.8%
GPT4-multi 6 9.8% 5 8.2% 4 9.8% 4 9.8% 10 9.8% 9 8.8%
PanicKiller 36 59.0% 27 44.3% 13 31.7% 11 26.8% 49 48.0% 38 37.3%
Regarding to the correct and plausible patches, we believe that, when coupled with the textual interpretations derived from our mined patterns, plausible patches can aid developers repair their program efficiently. This is because the failure of some patches to pass the regression tests may be attributed to additional logic after the patch is applied, which is beyond APR’s capabilities. Such plausible fixes, with minimal manual oversight, can be comprehensively resolved, thereby reducing the developers’ burden in understanding panic bugs and evaluating alternative functions.
5.3. Answer to RQ3: Efficiency
Fig 3 and Table 8 show the distribution of time spent on PanicKiller and manual fixes. On average, PanicKiller requires about one minute to repair panics, whereas manual repairs can take days. An interesting observation is that the average time of both PanicKiller and manual repair are generally longer for smaller datasets. For PanicKiller, this extended time is mainly due to the large test suites in some libraries, which lengthen the validation process. In terms of manual repairs, one notable case involved a panic bug that took over two years to fix, largely due to delayed maintenance by developers, resulting in a notably prolonged repair time for a smaller dataset.
Figure 3. Experiment results of efficiency.
In examining real-world PRs, we found a few key factors that potentially affect the bug-fixing time. Developers often spend several days discussing issues due to the need to comprehend error messages, understand collaborators’ code, and locate faults. Conversely, PanicKiller provides detailed explanations with fix patches to streamline the review process and improve developer understanding. Additionally, maintainers sometimes request multiple revisions of a submitted fix, requiring different approaches and thus increasing time and effort. PanicKiller addresses this by generating a range of patches, ranking them based on confidence scores and validation results, thus offering a prioritized patch list.
Table 8. Time consuming for PanicKiller and actual fixes.
Datasets Approaches Min Max Avg.
Panic4R- PanicKiller 0.51s 588.72s 65.77s
Small Actual 0.52h 807 days 76 days
Panic4R- PanicKiller 0.34s 425.32s 61.29s
Large Actual 0.42h 236 days 44 days
5.4. Answer to RQ4: Case Study
The overall evaluation results on real-world projects are presented in Table 9, showing that PanicKiller successfully resolved 28 of 41 open issues with varied root causes, with all patches merged by developers. As for unsuccessful fixes, some patches were plausible; they removed the panic but required further manual inspection and minor adjustments. In terms of closed issues, PanicKiller successfully generates 9 out of 22 patches that are similar to official patches. The proportion is not very high, largely due to differences in their repair locations. For example, while the official approach might add a branch before the parameter call, PanicKiller tends to insert it after creating the parameter. Nonetheless, both strategies achieve equivalent outcomes.
Table 9. Distribution of panics in real-world crates and fixing effectiveness of PanicKiller.
Crates Stars LoC Open issues Closed issues
Total Confirmed Total Similar
hifitime 315 10,762 28 21 0 -
unicode1 556 7,345 11 7 1 0
ratatui 9,086 41,534 1 0 18 7
fancy-regex 409 5,557 1 0 3 2
Total 41 28 22 9
  • 1
    The full name of this crate is: unicode-segmentation.
Below we illustrate some bug cases and their fixes.
Bug Case1: Listing 9 shows the resolution of an arithmetic overflow error, which is a proposed patch of a closed issue for Rust crate ratatui (ratatui:online, ). Through fault localization, PanicKiller succeeded in pinpointing a list of potential fault locations, prioritizing src/layout.rs:233:20 as the top candidate, which aligns exactly with the actual fix. Then, PanicKiller successfully generated 3 patches for this location. In this case, the fault was identified within a struct filed expression; however, a closer inspection through iterative analysis revealed that this match expression included a binary expression, making it suitable for applying the Mutate Binary Operator pattern. Finally, we calculate the similarity score between the generated patch and the official patch, obtaining a score of $0.95$, indicating the correctness.
{listing}
[h] {minted}[linenos=false, fontsize=]diff - x: self.x + margin.horizontal, + x: 0 + self.x.saturating_add(margin.horizontal), y: self.y + margin.vertical, A generated patch for the crate ratatui that fixes a panic bug caused by arithmetic overflow.
{listing}
[h] {minted}[linenos=false, fontsize=]diff - let mut cur_item = self.items[cur_item_idx].unwrap(); + let mut cur_item = match self.items[cur_item_idx] + Some(item) =¿ item, + None =¿ return Err(Errors::ParseError + (ParsingErrors::UnknownFormat)), + ; A generated patch that fixes a panic bug in hifitime.
Bug Case 2: The patch illustrated in Listing 9, generated by PanicKiller, serves as an example of addressing a panic bug that arises from unwrapping a None/invalid value. This bug comes from an open issue of Rust crate hifitime (hifitime:online, ), and our patch has been merged by developers. PanicKiller employs the Insert Match Unwrapper pattern, specifically designed to mitigate problems stemming from the misuse of unwrap. This approach transitions the unsafe usage of unwarp() to pattern matching, which results in returning an error message rather than triggering a panic error. Additionally, the patch created by PanicKiller takes into consideration the variable types to ensure consistency with the code context, thereby guaranteeing that the modified program passes regression testing.
6. Discussion
This section discusses our infrastructure’s application scenarios, comparison with existing works, and threats to validity.
6.1. Application Scenarios
This work establishes a systematic infrastructure that offers a diverse dataset and comprehensive analysis of panic bugs specific to Rust projects. We now briefly discuss the application scenarios and the potential future research directions.
  • Benchmark dataset for the Rust APR tools. Panic4R is designed to serve as a robust and reliable benchmark dataset specifically tailored for the evaluation of APR tools targeting Rust panic bugs. It includes macro switches and test case execution scripts, facilitating researchers to employ it as a benchmark dataset in their evaluations.
  • Extendable dataset for Rust program repair. The project structure and test scripts included in Panic4R are both unified and highly extensible. Researchers interested in Rust and program repair can leverage this foundational dataset to expand into other bug types or develop new repair patches.
  • Templates for pattern-based APR tools development. The fix patterns can serve as templates for pattern-based APR tools. Based on this foundation, researchers can concentrate on developing improved techniques for fault localization and pattern prioritization. In this paper, we develop PanicKiller as an application to prove the effectiveness of these patterns.
  • Dataset for fine-tuning large language models (LLMs). As LLMs rapidly evolve, using high-quality data for fine-tuning can enhance their domain-specific applications. By leveraging our carefully curated Panic4R, which includes natural language interpretations and code change examples, researchers could fine-tune LLMs to develop AI-driven techniques for panic bug repair.
  • Providing informative suggestions for users. PanicKiller produces a mix of suggestions—code patches and textual interpretations—that are more informative than compiler raw output. Given the steep learning curve of programs like Rust, we believe providing enhanced feedback is essential.
  • Helpful for fixing real-world opening panic bugs. Our experiment results have shown PanicKiller is much more efficient than manual fixes. Additionally, PanicKiller has successfully resolved 28 panic bugs in real Rust crates, demonstrating its practical effectiveness.
6.2. Comparison with Existing Code Patterns
Comparison with Java/C/C++ code patterns. Although numerous bug-fixing patterns oriented towards Java, C, and C++ have been proposed to corresponding APR tools, the code patterns we have identified are specific to Rust and are utilized for fixing panic bugs. Specifically, some of the patterns, e.g., Reorder State Changer, are designed to handle concurrency panics unique to Rust. The patterns such as Insert Unsafe Block are employed to address the unique unsafe features in Rust. Additionally, unlike Java or C++, where safety rules such as ownership are not enforced by the language, our proposed patterns carefully maintain these safety rules. This attention to Rust’s unique ownership model ensures that the fixes not only resolve the bugs but also uphold the language’s guarantees of memory safety and concurrency.
Comparison with Rust-specific code patterns. To date, only a few fix patterns have been specifically developed for Rust programs. Rust-lancet (yang2024lancet, ) and Ruxanne (robati2024patterns, ) are the primary studies that have designed bug and fix code patterns uniquely for Rust. Unlike these studies, which focus on ownership-related or other common bug types such as missing attributes, our infrastructure addresses the most critical panic bugs in Rust. While bugs violating lifetime rules and other common issues usually prevent compilation, offering error messages and corrective suggestions, panic bugs manifest at runtime, causing abrupt program termination. These bugs, influenced by Rust’s distinct memory and process management, render existing fix patterns inadequate for addressing panic issues. This paper introduces the first infrastructure aimed at exploring Rust panic bugs, identifying their root causes and developing specific repair patterns, significantly enhancing the understanding and remediation of Rust programs.
6.3. Threats to Validity
One potential threat to validity concerns the representativeness of our collected dataset and mined patterns, since all code and patches are sourced from open-source crates. However, we consider these crates to be relatively complex and representative of large-scale Rust projects. Panic4R comprises the top 500 most downloaded crates, reflecting the actual usage frequency and activity levels of these programs. We have also modeled our data collection process after the Defects4J dataset to enhance the reliability of our data. This adherence to proven methodologies in dataset construction supports the validity of our research findings.
7. Related Work
Automated Program Repair. Automated Program Repair (APR) has witnessed significant advancements in recent years. Most of methods (liu2019avatar, ; koyuncu2019ifixr, ; xia2022alapha, ; wen2018capgen, ; chen2017jaid, ; chen2021jaid, ; saha2019herclues, ; wong2021varfix, ; liu2019kpar, ; liu2019tbar, ) are designed for Java programs, with evaluated on Defect4J (just2014defect4j, ), a dataset of real bugs from open source Java programs. A recent study (huang2023survey, ) divides non-learning-based APR into three categories: search-based (weimer2013ae, ; qi2013trpautorepair, ; wen2018capgen, ; sidiroglou2015cp, ; le2012genprog, ), constraint-based (pei2011autofix, ; pei2014autofix, ; wei2010autofix, ; chen2017jaid, ; chen2021jaid, ; afzal2021sosrepair, ; mechtaev2015directfix, ) and template-based (tan2015relifix, ; le2016hdrepair, ; saha2019herclues, ). VarFix (wong2021varfix, ), a search-based way of observing which combinations of edit operations pass the test. Constraint-based techniques like Nopol (xuan2017nopol, ) and SemFix (nguyen2013semfix, ) transform the repair process into a constraint solving problem, reducing the search space. kPAR (liu2019kpar, ) and AVATAR (liu2019avatar, ) generate fix patterns collected by manual extraction and static violation analysis respectively, used by iFixR (koyuncu2019ifixr, ). TBar (liu2019tbar, ) is proposed to assess the qualitative and quantitative diversity of previous repair templates. As for learning-based APR, AlaphaRepair (xia2022alapha, ) achieves state-of-the-art results on both Java and Python programs via zero-shot learning. VulRepair (fu2022vulrepair, ) highlights the advancement of NMT-based automated vulnerability repairs with pre-trained models.
Different from existing APR tools that focus on Java and C programs (huang2023survey, ), we represents the first dedicated infrastructure targeted at Rust panic bugs, including a real-world dataset Panic4R and an APR tool PanicKiller designed for practical application. Several tools (yang2024lancet, ; robati2024patterns, ; deligiannis2023rustassistant, ) have been proposed to address Rust compilation bugs. However, they do not specifically target panic bugs that occur during runtime. Considering the significant impact of Defects4J and the steep learning curve associated with Rust, we believe Panic4R would serve as a foundational resource for Rust’s research.
Rust Program Testing and Analysing. Due to Rust’s innovative safety mechanism, new challenges have been posed in its testing and analyzing. RustSmith (sharma2023rustsmith, ) employs random program generation to test the Rust compiler. As for RULF (jiang2022rulf, ) and SyRust (takashima2021syrust, ), they concentrate on testing Rust crates by generating API sequences. In the realm of Rust program analysis, RUPTA (li2024rupta, ) introduces a context-sensitive pointer analysis framework for Rust, successfully applied to construct call graphs. Additionally, through static analysis, tools like SafeDrop (cui2023safedrop, ) and Rudra (bae2021rudra, ) detect memory safety issues in large-scale Rust programs, contributing to enhanced program robustness. In contrast, RustCheck (xia2023rustcheck, ) employs dynamic analysis techniques to uncover memory safety vulnerabilities.
Different from existing testing approaches, our work proposed the first APR tool specifically tailored to address errors related to Rust’s panic mechanism, and we focused on fixing the practical panic bugs. Besides, we have constructed a real-world code dataset and fix patterns, which serves as an infrastructure for Rust program comprehension and repair.
8. Conclusion
In this paper, we introduce an infrastructure designed to address Rust panic bugs. We construct the first Rust program’s fixing dataset, Panic4R, containing $102$ real-world panic bugs and their patches. Additionally, we conduct pattern mining to identify Rust-specific fixing patterns. We also introduce an APR tool, PanicKiller, which effectively localizes faults and generates patches, outperforming commercial LLM-based tools. Moreover, PanicKiller has successfully resolved 28 open issues related to panics, all of which have been confirmed and merged by developers.